OpenStudy (anonymous):

To rent a certain meeting room, a college charges a reservation fee of $38 and an additional fee of $8.70 per hour. The math club wants to spend less than $90.20 on renting the meeting room. What are the possible amounts of time for which they could rent the meeting room? Use t for the number of hours the meeting room is rented, and solve your inequality for t. (using >,=>,<,<=)

OpenStudy (anonymous):

I believe the inequality would be t => 6

OpenStudy (texaschic101):

38 + 8.7t <= 90.20 8.7t <= 90.20 - 38 8.7t <= 52.20 t <= 52.20/8.7 t < = 6

OpenStudy (anonymous):

but how would it be <. I thought => means less than or equal to. Meaning they could have the room for up to six hours

OpenStudy (texaschic101):

< means less then > means greater then picture you inequality sign as an alligators mouth opening....and eating the bigger number. Examples : 4 > 2......the alligators mouth is going after the bigger number...so 4 is greater then 2. 2 < 3....alligator is going after the 3 because it is bigger....therefore, 2 is less then 3. and = just means equals

OpenStudy (texaschic101):

your answer would be 6 hrs or less... so t <= 6